# DOS: Directional Object Separation in Text Embeddings for Multi-Object Image Generation

### v3

@@ 6 Limitation
We demonstrated that the adaptive strengths (defined in Eq. 5) in DOS effectively adjust the magnitude of separation vectors, leading to improved multi-object image generation (see Table 5). However, as shown in Figure 6, it does not always yield the optimal strengths. In this figure, we compare the default adaptive strength with the optimal strength obtained via grid search with a 0.1 interval. Although such cases are rare, they suggest that there is still room to improve how the strength of each separation vector is determined, potentially by incorporating additional criteria beyond shape, texture, and background bias.
@@ 7 Conclusion
DOS: Directional Object Separation in Text Embeddings
for Multi-Object Image Generation
@@ Appendix A Details on Preliminary Analysis
Aspect Condition Prompt Template Object Tuples
Shape Dissimilar
“a/an {objectAobject~A} and
a/an {objectBobject~B}”
(coin, brick), (donut, chocolate bar), (plate, ice cube), (basketball, television), (balloon, window), (coin, chocolate bar), (donut, soap bar), (basketball, picture frame), (balloon, monitor), (orange, sugar cube), (shoebox, clock), (soap bar, ring), (sugar cube, frisbee), (monitor, orange), (picture frame, bubble), (brick, ring), (ice cube, frisbee), (television, bubble), (picture frame, orange), (brick, bubble)
Similar
“a/an {objectAobject~A} and
a/an {objectBobject~B}”
(coin, clock), (donut, ring), (plate, frisbee), (basketball, orange), (balloon, bubble), (coin, ring), (donut, frisbee), (basketball, bubble), (balloon, orange), (orange, bubble), (shoebox, brick), (soap bar, chocolate bar), (sugar cube, ice cube), (monitor, television), (picture frame, window), (brick, chocolate bar), (ice cube, soap bar), (television, picture frame), (picture frame, monitor), (brick, sugar cube)
Texture Dissimilar
“a/an {objectAobject~A} and
a/an {objectBobject~B}”
(fur coat, coin), (peach, saucepan), (sweater, wrench), (cardigan, belt buckle), (rabbit, kettle), (fur coat, key), (mitten, knife), (sweater, saucepan), (puppy, kettle), (cardigan, wrench), (screw, moss patch), (key, mink coat), (bolt, mitten), (ring, towel), (medal, puppy), (screw, peach), (belt buckle, mink coat), (wrench, moss patch), (ring, peach), (saucepan, sweater)
Similar
“a/an {objectAobject~A} and
a/an {objectBobject~B}”
(fur coat, moss patch), (peach, mink coat), (sweater, mitten), (cardigan, towel), (rabbit, puppy), (fur coat, peach), (mitten, mink coat), (sweater, towel), (puppy, moss patch), (cardigan, sweater), (screw, coin), (key, saucepan), (bolt, wrench), (ring, belt buckle), (medal, kettle), (screw, key), (belt buckle, saucepan), (wrench, screw), (ring, kettle), (saucepan, wrench)
Background Bias Similar
“a/an {objectAobject~A} and
a/an {objectBobject~B}”
(goat, sheep), (monkey, chicken), (lion, deer), (pig, chicken), (cow, monkey), (chicken, lion), (goat, rabbit), (rabbit, gorilla), (chicken, sheep), (monkey, deer), (seahorse, octopus), (shrimp, whale), (dolphin, starfish), (jellyfish, shark), (shark, clownfish), (dolphin, seahorse), (starfish, whale), (seahorse, clownfish), (shrimp, shark), (tuna, shrimp)
Dissimilar
“a/an {objectAobject~A} and
a/an {objectBobject~B}”
(goat, octopus), (monkey, whale), (lion, starfish), (pig, shark), (cow, octopus), (chicken, seahorse), (goat, whale), (rabbit, clownfish), (chicken, shark), (monkey, shrimp), (seahorse, sheep), (shrimp, chicken), (dolphin, deer), (jellyfish, chicken), (shark, monkey), (dolphin, lion), (starfish, rabbit), (seahorse, gorilla), (shrimp, sheep), (tuna, deer)
Object Count Few
“a/an {objectAobject~A}”
(dog), (cat), (lion), (horse), (penguin), (rabbit), (gorilla), (bear), (cow), (monkey)
“a/an {objectAobject~A} and
a/an {objectBobject~B}”
(dog, cat), (cat, lion), (horse, penguin), (rabbit, gorilla), (penguin, horse), (bear, cow), (horse, lion), (rabbit, cat), (cow, penguin), (dog, lion)
Many
“a/an {objectAobject~A},
a/an {objectBobject~B}, and
a/an {objectCobject~C}”
(dog, cat, horse), (cat, lion, penguin), (horse, penguin, cat), (rabbit, gorilla, cow), (penguin, horse, gorilla), (bear, cow, horse), (horse, lion, cat), (rabbit, cat, dog), (cow, penguin, dog), (dog, lion, cat)
“a/an {objectAobject~A},
a/an {objectBobject~B},
a/an {objectCobject~C}, and
a/an {objectDobject~D}”
(dog, cat, horse, gorilla), (cat, lion, penguin, bear), (horse, penguin, cat, lion), (rabbit, gorilla, cow, cat), (penguin, horse, gorilla, rabbit), (bear, cow, horse, penguin), (horse, lion, cat, penguin), (rabbit, cat, dog, cow), (cow, penguin, dog, gorilla), (dog, lion, cat, rabbit)
@@ B.1 Implementation details
All experiments are conducted on a single NVIDIA RTX 3090 GPU. Our method builds on SDXL (PodellEtAl2023) and is also implemented with the recently released Stable Diffusion 3.5 (SD3.5) (stabilityai2024stable35). To construct the adaptive strengths defined in Eq. 5, we use 42 representative words covering a diverse range of object shapes and textures and 36 representative background phrases, which are selected through interaction with GPT-o4-mini-high (openai_gpt-o4-mini-high_2025). The full lists of these words and phrases are shown in Table 7. Table 8 presents the offsets used in the shifted tempered sigmoid function in Eq. 5, provided for each base model and embedding type. All offsets are precomputed over all object pairs from 80 MS-COCO (lin2014microsoft) classes.
@@ B.2 Benchmark details
We construct four benchmarks corresponding to the problematic scenarios reported in Figure 1, which are used in our main experiments: Similar Shapes, Similar Textures, and Dissimilar Background Biases, and Many Objects. Each of the first three benchmarks consists of 50 two-object prompts, following the template of “a/an $\{object~A\}$ and a/an $\{object~B\}$.” The object pairs for these benchmarks are selected through interactions with GPT-o4-mini-high (openai_gpt-o4-mini-high_2025). For example, to obtain 50 object pairs for the Similar Shapes benchmark, we ask the model: “Suggest 50 pairs of objects that share a similar overall shape or silhouette, even if they are unrelated in nature.” For the Many Objects benchmark, we extend the animal list used in Attend-and-Excite to include 17 animals, from which we randomly sample to create 25 three-object prompts and 25 four-object prompts. The extended animal list is defined as [cat, dog, bird, bear, lion, horse, elephant, monkey, frog, turtle, rabbit, fish, panda, cow, gorilla, penguin, chicken]. Table 9 shows the prompt templates and object tuples used to construct 50 prompts for each benchmark.
@@
Dissimilar
Background Biases
“a/an {objectAobject~A} and a/an {objectBobject~B}” (beach ball, snowball), (coconut, ice cube), (sandcastle, igloo), (kayak, snowboard), (camel, seahorse), (cow, whale), (goat, octopus), (chicken, fish), (chameleon, seahorse), (horse, seal), (chameleon, penguin), (rabbit, seal), (pineapple, coral), (apple, clam), (coconut, starfish), (lettuce, seaweed), (cauliflower, seahorse), (potato, oyster), (lion, whale), (camel, polar bear), (elephant, shark), (cactus, penguin), (penguin, camel), (fish, bicycle), (crocodile, eagle), (deer, shark), (rabbit, crab), (eagle, seahorse), (igloo, desert tent), (snowman, sandcastle), (ice skate, surfboard), (crocodile, camel), (koala, dolphin), (bear, crab), (cat, jellyfish), (giraffe, seahorse), (sofa, tent), (tractor, sailboat), (boat, car), (cactus, seahorse), (cactus, coral), (strawberry, fish), (parrot, penguin), (tank, canoe), (polar bear, kangaroo), (deer, sea turtle), (monkey, octopus), (carrot, seaweed), (carrot, coral), (dog, dolphin)
@@ B.3 Evaluation metrics
We adjust the VLM-based metric from EnMMDiT (wei2024enhancing) to measure both the success rate (SR) and mixture rate (MR) for each generated image. We provided GPT-4o-mini with the generated image and the text prompt used for image generation, asking it to classify each object in the prompt as (1) fully intact, (2) mixed, or (3) absent. SR is calculated as the ratio of images where all objects are classified as fully intact, and MR is calculated as the ratio of images where one or more objects are classified as mixed. Figure 7 shows examples illustrating our VLM-based evaluation.
@@ C.1 Combining DOS with Attend-and-Excite
Since DOS only modifies CLIP text embeddings, it can be used in combination with other latent modification methods such as Attend-and-Excite (A&E) (chefer2023attend). This combined method addresses some issues that cannot be solved by using either DOS or A&E alone, as shown in Figure 9. We further validate the effectiveness of this combination through a quantitative comparison, presented in Table 10.
Figure 9: Our method, DOS, can be used in combination with other latent modification method such as A&E (chefer2023attend). All methods are based on SDXL.
Table 10: The combination of DOS and A&E (A&E+Ours) demonstrates a synergistic effect, achieving best performance across all benchmarks.
Method
Similar
Shapes
Similar
Textures
Dissimilar
Background Biases
Many
Objects
@@ C.3 Extended Many Objects benchmark
To address concerns regarding the benchmark’s scope and to validate the external validity of our method, we constructed an extended Many Objects benchmark with greater categorical diversity. Moving beyond the original animal-only set, we incorporated objects from electronics, vehicles, and fruits. The expanded object list now consists of 16 distinct objects: [cat, dog, frog, fish, smartphone, laptop, television, camera, car, bicycle, motorcycle, train, apple, orange, avocado, strawberry]. From this set, we randomly generated 25 three-object and 25 four-object prompts. As shown in Table 11, the performance trends are consistent with those on the original benchmark presented in Table 1. This result suggests that our method’s effectiveness is not limited to a single category but generalizes across diverse object categories.
@@ C.4 Sensitivity to predefined words and phrases
Method
Similar
Shapes
Similar
Textures
Dissimilar
Background Biases
Many
Objects
@@ C.5 DOS vectors encode directional information
SR↑
MR↓
SR↑
(g-dino)
SR↑
MR↓
SR↑
(g-dino)
SR↑
MR↓
SR↑
(g-dino)
SR↑
MR↓
SR↑
(g-dino)
@@ C.6 Validation with multiple evaluators
While the Success Rate (SR) and Mixture Rate (MR) in Table 1 were measured using GPT-4o-mini (openai_gpt-4o-mini_2024), relying on a single Vision Language Model (VLM) could introduce evaluation biases. To validate the reliability of our evaluations, we re-evaluated all methods using two additional evaluators: another VLM, Gemini-2.0-Flash-Lite (google_gemini-2.0-flash-lite_2025), and an open-vocabulary object detector, Grounding-DINO (liu2024grounding). The MR metric was excluded for the object detector as it is not capable of assessing object mixing. The re-evaluated quantitative comparison, presented in Table 13, shows a consistent trend with our main results, suggesting their validity.
@@
Our main comparison in Table 1 was conducted on SDXL as the base model. To verify that this performance advantage extends to more recent architectures, we compared our method against TEBOpt (chen2024cat) and Self-Cross (qiu2025self) using SD3.5 as the base model. Table 14 shows that our approach maintains its superior performance on this newer architecture.
Table 14: Performance comparison on SD3.5 as the base model.
Method
Similar
Shapes
Similar
Textures
Dissimilar
Background Biases
Many
Objects
@@ C.8 Additional qualitative comparisons
Although our main benchmarks focus on failures arising from inter-object relationships in attribute‑free prompts, we additionally demonstrate that DOS maintains and can even improve attribute binding on the widely used Attend-and-Excite (A&E) benchmark (chefer2023attend). The benchmark defines three prompt templates: (i) “a {animalAanimal~A} and a {animalBanimal~B}”, (ii) “a {animalanimal} and a {colorcolor} {objectobject}”, and (iii) “a {colorAcolor~A} {objectAobject~A} and a {colorBcolor~B} {objectBobject~B}”, of which template (iii) is the most challenging because the model must correctly bind two independent color–object pairs in a single image. Figures 12 and 13 compare DOS with four baselines on template (iii): SDXL, TEBOpt, A&E, and CONFORM. Despite the absence of an explicit attribute–object binding guidance, DOS consistently keeps each color attached to the correct object and delivers clearer separation between the two objects, whereas competing methods often fail to bind attributes correctly or neglect the objects. Although the separation vectors in DOS were devised to mitigate inter‑object interference, they also implicitly preserve the binding between attributes and objects. These results suggest that DOS can be applied effectively to general prompts containing multiple attribute-object tuples, extending its utility beyond specific scenarios.

### v4

@@ 6 Limitation
We demonstrated that the adaptive strengths (defined in Eq. 5) in DOS effectively adjust the magnitude of separation vectors, leading to improved multi-object image generation (see Table 5). However, as shown in Figure 6, it does not always yield the optimal strengths. In this figure, we compare the default adaptive strength with the optimal strength obtained via grid search over 121 hyperparameter settings (an $\times11$ grid with a 0.1 interval). Although such cases are rare, they suggest that there is still room to improve how the strength of each separation vector is determined, potentially by incorporating additional criteria beyond shape, texture, and background bias.
@@ 7 Conclusion
Acknowledgments
This work was supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (No. RS-2022-NR070876) and in part by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) ([NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)], [No.RS-2023-00235293, Development of autonomous driving big data processing, management, search, and sharing interface technology to provide autonomous driving data according to the purpose of usage]).
@@ Appendix A Details on Preliminary Analysis
Aspect Condition Prompt Template Object Tuples
Shape Dissimilar “a/an {objectAobject~A} and a/an {objectBobject~B}” (coin, brick), (donut, chocolate bar), (plate, ice cube), (basketball, television), (balloon, window), (coin, chocolate bar), (donut, soap bar), (basketball, picture frame), (balloon, monitor), (orange, sugar cube), (shoebox, clock), (soap bar, ring), (sugar cube, frisbee), (monitor, orange), (picture frame, bubble), (brick, ring), (ice cube, frisbee), (television, bubble), (picture frame, orange), (brick, bubble)
Similar “a/an {objectAobject~A} and a/an {objectBobject~B}” (coin, clock), (donut, ring), (plate, frisbee), (basketball, orange), (balloon, bubble), (coin, ring), (donut, frisbee), (basketball, bubble), (balloon, orange), (orange, bubble), (shoebox, brick), (soap bar, chocolate bar), (sugar cube, ice cube), (monitor, television), (picture frame, window), (brick, chocolate bar), (ice cube, soap bar), (television, picture frame), (picture frame, monitor), (brick, sugar cube)
Texture Dissimilar “a/an {objectAobject~A} and a/an {objectBobject~B}” (fur coat, coin), (peach, saucepan), (sweater, wrench), (cardigan, belt buckle), (rabbit, kettle), (fur coat, key), (mitten, knife), (sweater, saucepan), (puppy, kettle), (cardigan, wrench), (screw, moss patch), (key, mink coat), (bolt, mitten), (ring, towel), (medal, puppy), (screw, peach), (belt buckle, mink coat), (wrench, moss patch), (ring, peach), (saucepan, sweater)
Similar “a/an {objectAobject~A} and a/an {objectBobject~B}” (fur coat, moss patch), (peach, mink coat), (sweater, mitten), (cardigan, towel), (rabbit, puppy), (fur coat, peach), (mitten, mink coat), (sweater, towel), (puppy, moss patch), (cardigan, sweater), (screw, coin), (key, saucepan), (bolt, wrench), (ring, belt buckle), (medal, kettle), (screw, key), (belt buckle, saucepan), (wrench, screw), (ring, kettle), (saucepan, wrench)
Background Bias Similar “a/an {objectAobject~A} and a/an {objectBobject~B}” (goat, sheep), (monkey, chicken), (lion, deer), (pig, chicken), (cow, monkey), (chicken, lion), (goat, rabbit), (rabbit, gorilla), (chicken, sheep), (monkey, deer), (seahorse, octopus), (shrimp, whale), (dolphin, starfish), (jellyfish, shark), (shark, clownfish), (dolphin, seahorse), (starfish, whale), (seahorse, clownfish), (shrimp, shark), (tuna, shrimp)
Dissimilar “a/an {objectAobject~A} and a/an {objectBobject~B}” (goat, octopus), (monkey, whale), (lion, starfish), (pig, shark), (cow, octopus), (chicken, seahorse), (goat, whale), (rabbit, clownfish), (chicken, shark), (monkey, shrimp), (seahorse, sheep), (shrimp, chicken), (dolphin, deer), (jellyfish, chicken), (shark, monkey), (dolphin, lion), (starfish, rabbit), (seahorse, gorilla), (shrimp, sheep), (tuna, deer)
Object Count Few “a/an {objectAobject~A}” (dog), (cat), (lion), (horse), (penguin), (rabbit), (gorilla), (bear), (cow), (monkey)
“a/an {objectAobject~A} and a/an {objectBobject~B}” (dog, cat), (cat, lion), (horse, penguin), (rabbit, gorilla), (penguin, horse), (bear, cow), (horse, lion), (rabbit, cat), (cow, penguin), (dog, lion)
Many “a/an {objectAobject~A}, a/an {objectBobject~B}, and a/an {objectCobject~C}” (dog, cat, horse), (cat, lion, penguin), (horse, penguin, cat), (rabbit, gorilla, cow), (penguin, horse, gorilla), (bear, cow, horse), (horse, lion, cat), (rabbit, cat, dog), (cow, penguin, dog), (dog, lion, cat)
“a/an {objectAobject~A}, a/an {objectBobject~B}, a/an {objectCobject~C}, and a/an {objectDobject~D}” (dog, cat, horse, gorilla), (cat, lion, penguin, bear), (horse, penguin, cat, lion), (rabbit, gorilla, cow, cat), (penguin, horse, gorilla, rabbit), (bear, cow, horse, penguin), (horse, lion, cat, penguin), (rabbit, cat, dog, cow), (cow, penguin, dog, gorilla), (dog, lion, cat, rabbit)
@@ B.1 Implementation details
All experiments are conducted on a single NVIDIA RTX 3090 GPU. Our method builds on SDXL (Podell et al. 2023) and is also implemented with the recently released Stable Diffusion 3.5 (SD3.5) (Stability AI 2024). To construct the adaptive strengths defined in Eq. 5, we use 42 representative words covering a diverse range of object shapes and textures and 36 representative background phrases, which are selected through interaction with GPT-o4-mini-high (OpenAI 2025). The full lists of these words and phrases are shown in Table 7. Table 8 presents the offsets used in the shifted tempered sigmoid function in Eq. 5, provided for each base model and embedding type. All offsets are precomputed over all object pairs from 80 MS-COCO (Lin et al. 2014) classes.
@@ B.2 Benchmark details
We construct four benchmarks corresponding to the problematic scenarios reported in Figure 1, which are used in our main experiments: Similar Shapes, Similar Textures, and Dissimilar Background Biases, and Many Objects. Each of the first three benchmarks consists of 50 two-object prompts, following the template of “a/an $\{object~A\}$ and a/an $\{object~B\}$.” The object pairs for these benchmarks are selected through interactions with GPT-o4-mini-high (OpenAI 2025). For example, to obtain 50 object pairs for the Similar Shapes benchmark, we ask the model: “Suggest 50 pairs of objects that share a similar overall shape or silhouette, even if they are unrelated in nature.” For the Many Objects benchmark, we extend the animal list used in Attend-and-Excite to include 17 animals, from which we randomly sample to create 25 three-object prompts and 25 four-object prompts. The extended animal list is defined as [cat, dog, bird, bear, lion, horse, elephant, monkey, frog, turtle, rabbit, fish, panda, cow, gorilla, penguin, chicken]. Table 9 shows the prompt templates and object tuples used to construct 50 prompts for each benchmark.
@@
Dissimilar Background Biases “a/an {objectAobject~A} and a/an {objectBobject~B}” (beach ball, snowball), (coconut, ice cube), (sandcastle, igloo), (kayak, snowboard), (camel, seahorse), (cow, whale), (goat, octopus), (chicken, fish), (chameleon, seahorse), (horse, seal), (chameleon, penguin), (rabbit, seal), (pineapple, coral), (apple, clam), (coconut, starfish), (lettuce, seaweed), (cauliflower, seahorse), (potato, oyster), (lion, whale), (camel, polar bear), (elephant, shark), (cactus, penguin), (penguin, camel), (fish, bicycle), (crocodile, eagle), (deer, shark), (rabbit, crab), (eagle, seahorse), (igloo, desert tent), (snowman, sandcastle), (ice skate, surfboard), (crocodile, camel), (koala, dolphin), (bear, crab), (cat, jellyfish), (giraffe, seahorse), (sofa, tent), (tractor, sailboat), (boat, car), (cactus, seahorse), (cactus, coral), (strawberry, fish), (parrot, penguin), (tank, canoe), (polar bear, kangaroo), (deer, sea turtle), (monkey, octopus), (carrot, seaweed), (carrot, coral), (dog, dolphin)
@@ B.3 Evaluation metrics
We adjust the VLM-based metric from EnMMDiT (Wei et al. 2024) to measure both the success rate (SR) and mixture rate (MR) for each generated image. We provided GPT-4o-mini with the generated image and the text prompt used for image generation, asking it to classify each object in the prompt as (1) fully intact, (2) mixed, or (3) absent. SR is calculated as the ratio of images where all objects are classified as fully intact, and MR is calculated as the ratio of images where one or more objects are classified as mixed. Figure 7 shows examples illustrating our VLM-based evaluation.
@@ C.1 Combining DOS with Attend-and-Excite
Since DOS only modifies CLIP text embeddings, it can be used in combination with other latent modification methods such as Attend-and-Excite (A&E) (Chefer et al. 2023). This combined method addresses some issues that cannot be solved by using either DOS or A&E alone, as shown in Figure 9. We further validate the effectiveness of this combination through a quantitative comparison, presented in Table 10.
Figure 9: Our method, DOS, can be used in combination with other latent modification method such as A&E (Chefer et al. 2023). All methods are based on SDXL.
Method Similar Shapes Similar Textures Dissimilar Background Biases Many Objects
@@
Table 10: The combination of DOS and A&E (A&E+Ours) demonstrates a synergistic effect, achieving the best performance across all benchmarks.
@@ C.3 Extended Many Objects benchmark
To address concerns regarding the benchmark’s scope and to validate the external validity of our method, we construct an extended Many Objects benchmark with greater categorical diversity. Moving beyond the original animal-only set, we incorporate objects from electronics, vehicles, and fruits. The expanded object list now consists of 16 distinct objects: [cat, dog, frog, fish, smartphone, laptop, television, camera, car, bicycle, motorcycle, train, apple, orange, avocado, strawberry]. From this set, we randomly generate 25 three-object and 25 four-object prompts. As shown in Table 11, the performance trends are consistent with those on the original benchmark presented in Table 1. This result indicates that our method’s effectiveness is not limited to a single category but generalizes across diverse object categories.
@@ C.4 Sensitivity to predefined words and phrases
Method Similar Shapes Similar Textures Dissimilar Background Biases Many Objects
@@ C.5 DOS vectors encode directional information
SR↑ (gemini) MR↓ (gemini) SR↑ (g-dino) SR↑ (gemini) MR↓ (gemini) SR↑ (g-dino) SR↑ (gemini) MR↓ (gemini) SR↑ (g-dino) SR↑ (gemini) MR↓ (gemini) SR↑ (g-dino)
@@ C.6 Validation with multiple evaluators
While the Success Rate (SR) and Mixture Rate (MR) in Table 1 were measured using GPT-4o-mini (OpenAI 2024), relying on a single Vision Language Model (VLM) could introduce evaluation biases. To validate the reliability of our evaluation results, we re-evaluated all methods using two additional evaluators: another VLM, Gemini-2.0-Flash-Lite (Google 2025), and an open-vocabulary object detector, Grounding-DINO (Liu et al. 2024). The MR metric was excluded for the object detector as it is not capable of assessing object mixing. The re-evaluated quantitative comparison, presented in Table 13, shows a consistent trend with our main results, suggesting their validity.
@@
Our main comparison in Table 1 was conducted on SDXL as the base model. To verify that this performance advantage extends to more recent architectures, we compared our method against TEBOpt (Chen et al. 2024) and Self-Cross (Qiu et al. 2025) using SD3.5 as the base model. Table 14 shows that our approach maintains its superior performance on this newer architecture.
Method Similar Shapes Similar Textures Dissimilar Background Biases Many Objects
@@ C.8 Additional qualitative comparisons
Although our main benchmarks focus on failures arising from inter-object relationships in attribute‑free prompts, we additionally demonstrate that DOS maintains and can even improve attribute binding on the widely used Attend-and-Excite (A&E) benchmark (Chefer et al. 2023). The benchmark defines three prompt templates: (i) “a {animalAanimal~A} and a {animalBanimal~B}”, (ii) “a {animalanimal} and a {colorcolor} {objectobject}”, and (iii) “a {colorAcolor~A} {objectAobject~A} and a {colorBcolor~B} {objectBobject~B}”, of which template (iii) is the most challenging because the model must correctly bind two independent color–object pairs in a single image. Figures 12 and 13 compare DOS with four baselines on template (iii): SDXL, TEBOpt, A&E, and CONFORM. Despite the absence of an explicit attribute–object binding guidance, DOS consistently keeps each color attached to the correct object and delivers clearer separation between the two objects, whereas competing methods often fail to bind attributes correctly or neglect the objects. Although the separation vectors in DOS were devised to mitigate inter‑object interference, they also implicitly preserve the binding between attributes and objects. These results suggest that DOS can be applied effectively to general prompts containing multiple attribute-object tuples, extending its utility beyond specific scenarios.